OpenStudy (anonymous):

A washer and a dryer cost $609 combined. The washer costs $41 less than the dryer. What is the cost of the dryer?

OpenStudy (anonymous):

$568

OpenStudy (anonymous):

let x is the price of the dryer so the price of the washer is x-41 so we have x+x-41=609 2x=609+41 2x=650 x=650/2 x=325 the dryer cost $325

OpenStudy (anonymous):

@ktnguyen1 can u help me with 3 more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A local hamburger shop sold a combined total of 618 hamburgers and cheeseburgers on Monday. There were 68 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Monday?

OpenStudy (anonymous):

@ktnguyen1

OpenStudy (anonymous):

hambergur=(618-68)/2=275 chesseburger=275+68=343

OpenStudy (anonymous):

so the answer is 275 or 343?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

275 or 343????

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

the #of Hamburgers sold is 275 the #of Cheeseburgers sold is 343

OpenStudy (anonymous):

A total of 500 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was three times the number of adult tickets sold . How many adult tickets were sold?

OpenStudy (anonymous):

@ktnguyen1

OpenStudy (anonymous):

x is tickets for adults sold 3x is tickets for students sold we have x+3x=500 4x=500 x=500/4 x=125 adult tickets sold is 125

OpenStudy (anonymous):

Huilan is 15 years younger than Thomas. The sum of their ages is 31 . What is Thomas's age?

OpenStudy (anonymous):

@ktnguyen1

OpenStudy (anonymous):

Huilan ages=(31-15)/2=16/2=8 Huilan is 8 year old Thomas is 8+15=23 year old

OpenStudy (anonymous):

can u help me with some numbers problems? @ktnguyen1

OpenStudy (anonymous):

what are the #s problem?